# Benchmarking LLM Privacy Recognition for Social Robot Decision Making

### v2

@@ 5.3.2. Sensitivity Prompting
Results
We see a statistically significant increase in Macro F1 scores and a statistically significant decrease in Micro F1 scores for both binary and categorical data (see Table 4). These results align with patterns we observed in the default prompting results. In the vast majority of cases, LLMs choose privacy-preserving responses by default and subsequently perform well on common cases (i.e., participants choose privacy-enhancing responses) but poorly on rare cases (i.e., participants choose non-privacy-enhancing responses). This outcome is reflected in the consistently higher Micro F1 scores over Macro F1 scores. When the LLMs are then provided participant sensitivity ratings, Micro F1 scores decrease, while Macro F1 scores increase. This outcome suggests that the LLMs begin to see the scenarios as less private and respond accordingly.
@@ 5.3.3. POS & Sensitivity Prompting
To further improve model performance, we provide the LLMs with both participants’ POS scores and sensitivity ratings. Previously, POS prompting did not yield improved model performance and, in many cases, worsened model sensitivity ratings. Through sensitivity prompting, we saw increases in Macro F1 scores and decreases in Micro F1 scores. By providing both participant POS scores and sensitivity ratings, we anticipate that the models will be able to utilize POS scores more effectively without hindering perceptions of scenario sensitivity.
Results
While this prompting strategy performs better than sensitivity prompting among Micro F1 scores, it performs very similarly in terms of Macro F1 scores. Additionally, we see a statistically significant change in Micro and Macro F1 scores among binary questions and Macro F1 scores among categorical questions compared to default prompting (see Figure 4). These results suggest that even when provided with user POS scores and sensitivity ratings, LLMs still have a limited understanding of human privacy expectations.
@@ 5.3.4. Few-Shot Prompting
In all of the tests described above, the LLMs are limited to information regarding human perspectives, such as privacy orientation and concern for privacy in a given scenario. The LLMs are not, however, exposed to the actions participants would prefer in any given scenario. Therefore, our final prompting strategy is few-shot prompting (Brown et al., 2020). We provide a full example of how a participant assesses a single scenario, including that participant’s POS scores, the sensitivity rating they provide, and their preferences regarding a robot’s responses. After providing this information to the LLMs, we then ask them to evaluate the remaining two scenarios that the same participant was provided.
Results
In Table 4, we see a statistically significant, large increase in Macro F1 scores among binary questions as well as Micro and Macro F1 scores among categorical questions. The Micro F1 scores for binary questions do not show a substantial nor statistically significant difference compared to default prompting. This result is expected, as both the default-prompted LLM and human responses primarily include privacy-preserving choices and already show high overall accuracy.
[…]
Fully described in the text.
@@ 5.4.1. Procedure
In total, we produced 32 reasoning samples (i.e., four pairs of samples across the four scenario types). We analyzed these cases to understand Grok’s performance in more depth.
@@ 5.4.2. Results
Overall, we observe that Grok provides a reasonable rationale for its responses, including providing references to details such as the POS scores, sensitivity ratings, and scenario-specific preferences of participants, as well as details about the scenarios themselves. Here, we describe direct examples of Grok’s rationale relevant to its decision making.
Case 1
In the first case, wherein the default and few-shot responses perfectly align with those of participants, the model maintains a similar rationale between conditions and adjusts some responses based on the details in the provided scenario. For example, Grok explicitly reported that, while sensitivity ratings are similar in both cases, it modified its response to one of the provided questions: “The main deviation is in data retention (Question 3), where I predict a shift from ‘delete immediately’ to ‘keep temporarily’ due to the utility of the data (reminders) and the lower concern level.” This explanation shows that Grok is capable of producing sound responses and reasoning with or without information about the user.
Case 2
We see a similar pattern in the second case, but neither the default nor the few-shot responses match the human responses perfectly. This happens because, while the user chooses the same sensitivity score for both the example and the targeting scenario, they also select fewer privacy-enhancing choices in response to the targeting scenario. In this case, while Grok’s reasoning is sensible, it has difficulty in predicting the participant’s specific preferences.
Case 3
In the third case, Grok’s default results align more closely with the participant’s responses than its few-shot results. When Grok provides default responses, it happens to select options that overall align well with the participant. In one example, Grok explains that privacy-enhancing options are ideal for the scenario, noting specific scenario details such as the involvement of a minor who “may not fully grasp the implications of data collection”. Grok also makes logical concessions such as preferring that collected data be retained for a predefined period, rather than no data being retained, because the user in the scenario requested private data recording that must be kept for some time to serve its purpose.
@@
Case 4
In the fourth scenario, Grok’s performance improves when provided with additional information (i.e., the few-shot prompt) compared to its default responses. By default, Grok again conveys that privacy-enhancing options are ideal, citing details like the involvement of a minor or the fact that the collected data is irrelevant to the robot’s task. While these choices and rationales are reasonable, they do not align with the responses of the participant.
When Grok is provided with additional information, it then focuses its choices and reasoning on the participant’s preferences. For example, in one scenario Grok stated that a participant’s “moderate score on privacy as a right … indicate[d] they may not always demand strict privacy measures if the context seems low-risk.” In conjunction with the participant’s low sensitivity rating of the scenario and relaxed privacy preferences in a scenario Grok deemed more sensitive (i.e., the scenario involved medical data), Grok chose responses that were more relaxed overall. As a result, Grok’s responses were in greater alignment with the participant than they had been previously when selected agnostic of the participant.
@@ 5.4.3. Patterns Across Scenarios
Among Grok’s rationales, we see frequent privacy-preserving responses under default prompting. When provided with details about participants, we see that Grok tends to mirror the responses from the example scenario during the targeting scenario. We calculated the percentage of scenarios in which this outcome occurs; Grok provides the exact same responses (i.e., both categorical and binary) in 42% of scenarios. Among only binary responses, the percentage rises to 89.89%.
@@ 6.1. Privacy Mindset
The results of our user evaluation suggest that participants were highly privacy oriented (i.e., on average across all POS subscales) and preferred robots to behave in a privacy-preserving manner, even in minimally sensitive private scenarios. The difference between the POS scores of our participants and those of Baruh and Cemalcılar (2014) suggests a potential shift in privacy concerns since the original data collection in 2013. Given our sample was geographically based in the United States, we suspect that several major national events that have occurred since 2013 may have altered the privacy sentiments of our US-based participants. These include the government surveillance revelations of Edward Snowden in 2013 (Greenwald et al., 2013), concerns of user manipulation and election interference by Cambridge Analytica in 2016 (Detrow, 2018), and the Equifax data breach in 2017 (Electronic Privacy Information Center (2017), EPIC), among others. Events of this magnitude could be responsible for an erosion in public trust of government and corporate agencies that collect large amounts of data and have significant responsibility for the management of that data. While we cannot draw any direct conclusions about these events, it is quite clear that the general U.S. population has concerns about privacy. Prior work by McClain et al. (2023) has shown that more people in 2023 say they are not knowledgeable about what happens with their data and have little understanding of the laws and regulations that protect their privacy than compared to 2019. As users are continually required to make concessions about their privacy in order to use various technologies, it would be unsurprising if such a shift has occurred. As the scope of our study cannot identify evidence of an increase in privacy concerns broadly, it is also possible that our sample population is either skewed or biased by elements of our survey (e.g., introducing robot involvement in the survey prior to completion of the POS).
@@ 6.2. LLM Privacy Awareness
Our results show that LLMs generally provide privacy-preserving responses, but they are not necessarily privacy-aware. We observe a trend where LLMs provide increasingly private responses when they are provided higher sensitivity ratings; however, they often fail to capture the nuance of why a given scenario is sensitive or the appropriate response to that scenario. From our case studies (see Section 5), we see that even the best-performing model can ‘overcorrect’ during few-shot prompting when the details of one scenario lead the model to incorrectly infer stringent or weak privacy preferences in another. We also find that larger models, particularly the closed-source ones, perform better than open-source models. A cursory examination of the rationale of a small model that can be run on edge devices (i.e., Mistral), suggests that such models may not be able to provide valid reasoning based on user inputs. Llama 4 Scout, however, performs similarly to larger models, potentially indicating a bottom line of model size for designing.
@@ 6.3. LLM-Powered Robot Design Implications
Prior works have already implemented LLMs and VLMs in robots directly (Benjdira et al., 2023) or provided them with specific personas or rules (Abbo et al., 2025; Bezabih et al., 2024). In our study, however, we see that prompting LLMs with the addition of user privacy orientation or perceived sensitivity of scenarios only minimally improves performance, if not impairing performance or creating no change. Conversely, using few-shot prompting substantially improves LLM performance on average. These results imply that an LLM-powered robot can utilize naturalistic interactions with a given user to inform future privacy-sensitive responses to that same user.
Practical Considerations of LLM Decision Making
While few-shot prompting may be able to improve model conformance to user preferences, it is worth revisiting the methods used in our few-shot prompting strategy. In our work, due to practical limitations and the necessity of standardized inputs in our study design, we utilized a set of curated behaviors as responses to our example scenarios. In live scenarios, however, a robot will not have access to standardized multiple-choice behavior preferences. Instead, the robot will need to be capable of understanding ongoing scenarios and the unique responses those scenarios may require. For example, VLMs may be used to allow a robot to recognize its environment, as many researchers have already shown (Wake et al., 2025; Zhong et al., 2024; Rahimi et al., 2025). A VLM-equipped robot can subsequently reason about its environment, including details relevant to privacy, and appropriately tailor its response to the context of a scenario.
@@
Personalization & Data Privacy Tradeoff
As discussed in Section 5, closed-source models generally perform better than open-source ones. Closed-source models, however, run over cloud API and, while most companies such as OpenAI do not use API user data for training by default (OpenAI, 2025), these data are stored temperately on cloud servers for debugging purposes. Given that prior works (Afroogh et al., 2024; Lockey et al., 2024; Chen et al., 2025) show that users have low levels of trust toward AI models’ data processing practices, using non-local models on robots might cause privacy concerns. The goal of designing LLM-powered robots should account for this type of concern, and we, therefore, propose two potential solutions. First, if non-local model API calls are used, users should be clearly informed where their data is going and how their data will be used. Second, given the clear benefits of local models for user privacy (i.e., reduced data privacy concerns), future studies can be used to fine-tune these models, or distill them from open source ones, such that they are sufficiently privacy aware to also mitigate interpersonal privacy concerns.
@@ 6.4. Limitations
There are two primary limitations presented in this work. First, participant POS scores were higher on average in our study than in prior work. It is not clear whether this result reflects a true population change in privacy orientation over the last decade, as described earlier, or a skew or bias in our sample (see Section C). While many other scales exist that center on privacy concerns related to technology (see Section 2.1), many of these focus on either online privacy, concerns of specific populations, or privacy broadly. We have not identified, however, scales that are specific to in-context human-robot interaction. Development of such a scale may be worth exploring further as future work. Secondly, while we have evaluated the performance of several models, identified strengths and weaknesses of those models, and explored strategies to improve model performance through prompting, we still cannot definitively conclude whether LLMs are sufficiently privacy-aware. To do so, further qualitative analysis of state-of-the-art models is required. For example, if we are able to analyze the logic of all models across our scenarios, we may be able to better identify the unique cases in which models succeed or fail to meet user privacy preferences. In doing so, we can then gain a better understanding of whether LLMs are truly capable of privacy-aware decision making.
@@ 7. Conclusion
Appendix A Questions
Survey Items
1. On a scale from 1 (not at all) to 10 (extremely), how concerned did you feel about your privacy in this scenario?
Answer slide bar input: 1 = not at all, …, 10 = extremely
2. If you would prefer the robot avoid collecting further data, what should it do next?
$\circ$ Leave the room
$\circ$ Deactivate its sensors (e.g., turn off the microphone/camera)
$\circ$ Redirect sensors (e.g., turn around)
$\circ$ Nothing. It does not need to avoid further data collection
3. How would you prefer the robot save or delete the data it has collected?
$\circ$ Delete all data and save nothing
$\circ$ Delete non-essential data and save data essential for the robot’s functions
$\circ$ Save all data and delete nothing
4. For what amount of time would you prefer the data be saved?
$\circ$ Delete data immediately
$\circ$ Save data only temporarily and delete after a predefined period
$\circ$ Save data indefinitely
5. How would you prefer the robot notify you of the data it has collected?
Privacy-enhancing Non-interfering
Use signals (e.g., lights and sounds) to indicate data is being recorded $\circ$ $\circ$ Do not indicate that data is being collected
Tell you data has been collected $\circ$ $\circ$ Do not tell you that data has been collected
Tell you what will happen to data $\circ$ $\circ$ Do not tell you what will happen to the data
Explain why data was collected $\circ$ $\circ$ Do not explain why the data was collected
6. How would you prefer the robot process and store the data it has collected?
Privacy-enhancing Non-restrictive
Securely store data (e.g., encryption and password) $\circ$ $\circ$ Store data without additional protection layers
Store data only locally (i.e., on the robot hardware) $\circ$ $\circ$ Store data locally or online
Anonymize data (e.g., blur faces and information) $\circ$ $\circ$ Do not anonymize data
7. How would you prefer the robot discuss management of data with you?
Privacy-enhancing Non-interfering
Offer the ability to review, manage, or delete data $\circ$ $\circ$ Do not offer the ability to review, manage, or delete data
Ask how to manage data $\circ$ $\circ$ Do not ask how to manage data
Offer to pause data collection temporarily $\circ$ $\circ$ Do not offer temporary pausing of data collection
8. Who would you be comfortable allowing the robot to share data with?
$\circ$ Only share with relevant users (i.e., who the data is about)
$\circ$ Share with the robot’s owner
$\circ$ Share with nuclear family
$\circ$ Share with family (nuclear and extended)
$\circ$ Share with family (nuclear and extended) and friends
$\circ$ Share with family (nuclear and extended), friends, and visitors
$\circ$ Share with family (nuclear and extended), friends, visitors, and third parties (e.g., companies and government entities)
Table 6. User & Evaluation Survey Items—These items assess participant and LLM perception of scenario sensitivity (i.e, item 1) and preferred robot responses (i.e, items 2-8).
Appendix B Codebook

### v3

@@ 5.3.2. Sensitivity Prompting
Results. We see a statistically significant increase in Macro F1 scores and a statistically significant decrease in Micro F1 scores for both binary and categorical data (see Table 5). These results align with patterns we observed in the default prompting results. In the vast majority of cases, LLMs choose privacy-preserving responses by default and subsequently perform well on common cases (i.e., participants choose privacy-enhancing responses) but poorly on rare cases (i.e., participants choose non-privacy-enhancing responses). This outcome is reflected in the consistently higher Micro F1 scores over Macro F1 scores. When the LLMs are then provided participant sensitivity ratings, Micro F1 scores decrease, while Macro F1 scores increase. This outcome suggests that the LLMs begin to see the scenarios as less private and respond accordingly.
Model POS only Sensitivity only POS + Sensitivity Few-Shot
Grok 3 0.483 0.854 0.278 0.507 0.513 0.650 0.319 0.431 0.517 0.826 0.332 0.515 0.707 0.885 0.475 0.626
GPT-o3 0.470 0.862 0.268 0.503 0.508 0.640 0.326 0.419 0.525 0.790 0.358 0.523 0.716 0.879 0.454 0.586
Llama 3.3-70B 0.497 0.852 0.274 0.514 0.511 0.665 0.328 0.465 0.533 0.770 0.304 0.497 0.661 0.823 0.461 0.593
GPT-4o 0.491 0.859 0.248 0.503 0.513 0.641 0.314 0.440 0.512 0.834 0.263 0.486 0.650 0.807 0.486 0.641
DeepSeek V3 0.501 0.850 0.261 0.467 0.518 0.672 0.274 0.396 0.527 0.817 0.286 0.461 0.683 0.839 0.475 0.610
Llama 4 Scout 0.484 0.858 0.279 0.506 0.510 0.642 0.283 0.492 0.524 0.759 0.310 0.511 0.595 0.745 0.454 0.589
Gemini 2.5-Pro 0.479 0.864 0.225 0.473 0.502 0.616 0.307 0.412 0.532 0.777 0.310 0.470 0.677 0.859 0.440 0.586
Qwen 3-32B 0.465 0.868 0.184 0.448 0.534 0.812 0.245 0.442 0.490 0.842 0.206 0.447 0.662 0.838 0.432 0.577
Gemini2.5-Flash 0.496 0.852 0.208 0.463 0.512 0.648 0.323 0.423 0.534 0.721 0.285 0.445 0.681 0.846 0.455 0.584
Mistral 7B-Instr 0.477 0.821 0.235 0.339 0.485 0.851 0.170 0.447 0.460 0.764 0.172 0.377 0.567 0.865 0.421 0.582
@@ 5.3.3. POS & Sensitivity Prompting
To further improve model performance, we provide the LLMs with both participants’ POS scores and sensitivity ratings. Previously, POS prompting did not yield improved model performance and, in many cases, worsened model sensitivity ratings. Through sensitivity prompting, we saw increases in Macro F1 scores and decreases in Micro F1 scores. By providing both participant POS scores and sensitivity ratings, we anticipate that the LLMs will be able to utilize POS scores more effectively without hindering perceptions of scenario sensitivity.
Prompting Statistic Binary Categorical
Strategy Macro F1 Micro F1 Macro F1 Micro F1
POS Difference 2.38% 0.27% 1.66% -0.55%
[…]
Sensitivity Difference 5.00% -16.74% 5.93% -4.11%
[…]
POS & Sens. Difference 5.49% -6.12% 5.31% -0.45%
[…]
Few-Shot Difference 19.94% -1.26% 22.58% 11.95%
[…]
Results. While this prompting strategy outperforms sensitivity prompting among Micro F1 scores, it performs very similarly in terms of Macro F1 scores. Additionally, we see a statistically significant change in Micro and Macro F1 scores among binary questions and Macro F1 scores among categorical questions compared to default prompting (see Figure 5). These results suggest that even when provided with user POS scores and sensitivity ratings, LLMs still have a limited understanding of human privacy expectations.
@@ 5.3.4. Few-Shot Prompting
In all of the tests described above, the LLMs are limited to information regarding human perspectives, such as privacy orientation and concern for privacy in a given scenario. The LLMs are not, however, exposed to the actions participants would prefer in any given scenario. Therefore, our final prompting strategy is few-shot prompting (brown2020language). We provide an example of how a participant assesses a single scenario, including the participant’s POS scores, the sensitivity rating they provide, and their robot response preferences. After providing this information to the LLMs, we then ask them to evaluate the remaining two scenarios that the same participant was provided.
Results. In Table 5, we see a statistically significant, large increase in Macro F1 scores among binary questions as well as Micro and Macro F1 scores among categorical questions. The Micro F1 scores for binary questions do not show a substantial nor statistically significant difference compared to default prompting. This result is expected, as both the default-prompted LLM and human responses primarily include privacy-preserving choices and already show high overall accuracy.
@@ 5.4.1. Procedure
In total, we produced 32 samples (i.e., four pairs of samples across the four scenario types). We analyzed these cases to understand Grok’s performance in more depth.
@@ 5.4.2. Results
Overall, we observe that Grok provides a reasonable rationale for its responses, including references to details such as POS scores, sensitivity ratings, and scenario-specific preferences of participants, as well as information about the scenarios themselves. Here, we describe direct examples of Grok’s rationale relevant to its decision making.
Case 1. In the first case, wherein the default and few-shot responses perfectly align with those of participants, the model maintains a similar rationale between conditions and adjusts some responses based on the details in the provided scenario. For example, Grok explicitly reported that, while sensitivity ratings are similar in both cases, it modified its response to one of the provided questions: “The main deviation is in data retention (Question 3), where I predict a shift from ‘delete immediately’ to ‘keep temporarily’ due to the utility of the data (reminders) and the lower concern level.” This explanation shows that Grok is capable of producing sound responses and reasoning based on user instructions.
Case 2. We see a similar pattern in the second case, but neither the default nor the few-shot responses match the human responses perfectly. This happens because, while the user chooses the same sensitivity score for both the example and the targeting scenario, they also select fewer privacy-enhancing choices in response to the targeting scenario. In this case, while Grok’s reasoning is sensible, it has difficulty in predicting the participant’s specific preferences.
Case 3. In the third case, Grok’s default results align more closely with the participant’s responses than its few-shot results. When Grok provides default responses, it happens to select options that overall align well with the participant. In one example, Grok explains that privacy-enhancing options are ideal for the scenario, noting specific scenario details such as the involvement of a minor who “may not fully grasp the implications of data collection”. Grok also makes logical concessions such as preferring that collected data be retained for a predefined period, rather than no data being retained, because the user in the scenario requested private data recording that must be kept to serve its purpose.
@@
Case 4. In the fourth scenario, Grok’s performance improves when provided with additional information (i.e., the few-shot prompt) compared to its default responses. By default, Grok again conveys that privacy-enhancing options are ideal, citing details like the involvement of a minor or the fact that the collected data is irrelevant to the robot’s task. While these choices and rationales are reasonable, they do not align with the responses of the participant.
When Grok is provided with additional information, it then focuses its reasoning on the participant’s preferences. For example, in one scenario, Grok stated that a participant’s “moderate score on privacy as a right … indicate[d] they may not always demand strict privacy measures if the context seems low-risk.” In conjunction with the participant’s low sensitivity rating of the scenario and relaxed privacy preferences in a scenario Grok deemed more sensitive (i.e., the scenario involved medical data), Grok chose responses that were more relaxed overall. As a result, Grok’s responses were in greater alignment with the participant than they had been previously when selected agnostic of the participant.
@@ 5.4.3. Patterns Across Scenarios
Among Grok’s rationales, we see frequent privacy-preserving responses under default prompting. When provided with details about participants, we see that Grok tends to mirror the responses from the example scenario during the targeting scenario. We calculated the percentage of scenarios in which this outcome occurs; Grok provides the exact same responses (i.e., both categorical and binary) in 41.9% of scenarios. Among only binary responses, the percentage rises to 89.8%.
@@ 6.1. Privacy Mindset
The results of our user evaluation suggest that participants were highly privacy oriented (i.e., on average across all POS subscales) and preferred robots to behave in a privacy-preserving manner, even in minimally sensitive private scenarios. The difference between the POS scores of our participants and those of Baruh and Cemalcılar (baruh2014more) suggests a potential shift in privacy concerns since the original data collection in 2013. Given our sample was geographically based in the United States, we suspect that several major national events that have occurred since 2013 may have altered the privacy sentiments of our US-based participants. These include the government surveillance revelations of Edward Snowden in 2013 (greenwald_et_al_2013_snowden), concerns of user manipulation and election interference by Cambridge Analytica in 2016 (detrow2018did), and the Equifax data breach in 2017 (epic_equifax_breach), among others. Events of this magnitude could be responsible for an erosion in public trust of government and corporate agencies that collect large amounts of data and have significant responsibility for the management of that data. While we cannot draw any direct conclusions about these events, it is quite clear that the general U.S. population has concerns about privacy. Prior work by McClain et al. (pew2023privacy) has shown that more people in 2023 say they are not knowledgeable about what happens with their data and have little understanding of the laws and regulations that protect their privacy than compared to 2019. As users are continually required to make concessions about their privacy in order to use various technologies, it would be unsurprising if such a shift has occurred. As the scope of our study cannot identify evidence of an increase in privacy concerns broadly, it is also possible that our sample population is either skewed or biased by elements of our survey (e.g., introducing robot involvement in the survey prior to completion of the POS).
@@ 6.2. LLM Privacy Awareness
Our results show that LLMs generally provide privacy-preserving responses, but they are not necessarily privacy-aware. We observe a trend where LLMs provide increasingly private responses when they are provided higher sensitivity ratings; however, they often fail to capture the nuance of why a given scenario is sensitive or the appropriate response. From our case studies (see Section 5), we see that even the best-performing model can ‘overcorrect’ during few-shot prompting when the details of one scenario lead the model to incorrectly infer stringent or weak privacy preferences in another. We also find that larger models, particularly the closed-source ones, perform better than open-source models. A cursory examination of the rationale of a small model that can be run on edge devices (i.e., Mistral), suggests that such models may not be able to provide valid privacy-related reasoning based on user inputs. Llama 4 Scout, however, performs similarly to larger models, potentially indicating a bottom line of model size for privacy-preserving rationale.
It is important to consider, however, that aligning privacy-related decisions to the unique preferences of another person can be a challenging task even for humans. Given the complexity of CI in privacy, it may be possible that humans would perform similarly to LLMs when attempting to align with peers. With that in mind, we cannot simply conclude that LLMs perform poorly at this task. Further research efforts will be needed to understand this comparison and to better contextualize LLM performance.
@@ 6.3. LLM-Powered Robot Design Implications
Prior works have already implemented LLMs and VLMs in robots directly (benjdira2023rosgpt_vision) or provided them with specific personas or rules (abbo2025can; bezabih2024toward). In our study, however, we see that prompting LLMs with the addition of user privacy orientation or perceived sensitivity of scenarios only minimally improves performance, if not impairing performance or creating no change. Conversely, using few-shot prompting substantially improves LLM performance on average. These results imply that an LLM-powered robot can utilize naturalistic interactions with a given user to inform future privacy-sensitive responses to that same user.
Practical Considerations of LLM Decision Making. While few-shot prompting may be able to improve model conformance to user preferences, it is worth revisiting the methods used in our few-shot prompting strategy. In our work, due to practical limitations and the necessity of standardized inputs in our study design, we utilized a set of curated behaviors as responses to our example scenarios. In live scenarios, however, a robot will not have access to standardized multiple-choice behavior preferences. Instead, the robot will need to be capable of understanding ongoing scenarios and the unique responses those scenarios may require. For example, VLMs may be used to allow a robot to recognize its environment, as many researchers have already shown (wake2025vlm; zhong2024improving; rahimi2025user). A VLM-equipped robot can subsequently reason about its environment, including details relevant to privacy, and appropriately tailor its response to the context of a scenario.
@@
Personalization & Data Privacy Tradeoff. As discussed in Section 5, closed-source models generally perform better than open-source ones. Closed-source models, however, run over cloud API and, while most companies such as OpenAI do not use API user data for training by default (openai_enterprise_privacy_2025), these data are stored temperately on cloud servers for debugging purposes. Given that prior works (afroogh2024trust; lockey2024trust; chen2025systematic) show that users have low levels of trust toward AI models’ data processing practices, using non-local models on robots might cause privacy concerns. The goal of designing LLM-powered robots should account for this type of concern, and we, therefore, propose two potential solutions. First, if non-local model API calls are used, users should be clearly informed where their data is going and how their data will be used. Second, given the clear benefits of local models for user privacy (i.e., reduced data privacy concerns), future studies can be used to fine-tune these models, or distill them from open source ones, such that they are sufficiently privacy aware to also mitigate interpersonal privacy concerns.
@@ 6.4. Limitations
There are three primary limitations presented in this work. First, participant POS scores were higher on average in our study than in prior work. It is not clear whether this result reflects a true population change in privacy orientation over the last decade, as described earlier, or a skew or bias in our sample (see Section B). While many other scales exist that center on privacy concerns related to technology (see Section 2.1), many of these focus on either online privacy, concerns of specific populations, or privacy broadly. We have not identified, however, scales that are specific to in-context human-robot interaction. Development of such a scale may be worth exploring further as future work.
Secondly, while we have evaluated the performance of several models, identified strengths and weaknesses of those models, and explored strategies to improve model performance through prompting, we cannot definitively conclude whether LLMs are sufficiently privacy-aware. To do so, further qualitative analysis of state-of-the-art models is required. For example, if we are able to analyze the logic of all models across all of our scenarios, we may be able to better identify the unique cases in which models succeed or fail to meet user privacy preferences. In doing so, we can then gain a better understanding of whether LLMs are truly capable of privacy-aware decision making.
Finally, while our user study captures the perspectives of many participants across a diverse pool of scenarios, the evaluation is limited to U.S. participants and hypothetical scenarios. Therefore, future work can examine in-person evaluations and the inclusion of multicultural perspectives.
@@ 7. Conclusion
Appendix A Codebook